A plane makes intercepts of 1, 2 and 3 Å on the crystallographic axes of an orthorhombic crystal with

a:b:c = 3:2:1. Determine the Miller indices of this plane. 


1
Expert's answer
2021-01-23T15:53:09-0500

Answer

Firstly see intercepts are

1A,2 A,3A respective to X, Y, Z axis.---(1)

Lattice parameters are 3L,2L,1L---(2)

Where L is any constant.

Now devide 1 to 2

"\frac{1}{3L}:\frac{2}{2L}:\frac{3}{1L}"

Take reciprocal

"\frac{3L} {1} :\frac{2L} {2} \frac{1L}{3}"

Leave L

And L and make in integer form

"(h, k, l) =(6 \space3\space1)"